OpenStudy (anonymous):

(4(pir^3-600))/r^2=0 what is r?

OpenStudy (anonymous):

\[\frac{ 4(\pi r^3-600) }{ r^2 }=0\]

OpenStudy (anonymous):

(4(pir^3-600))/r^2=0 (pir^3-600)=0 r^3=600/pi r= cube root(600/pi) ans.. or r=(600/pi)^1/3 ans.. use your calculator to simplify them more

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

ok YW have fun now and good luck ... :D